### v5

# On Bessel’s Correction: Unbiased Sample Variance, the Bariance, and a Novel Runtime-Optimized Estimator

Abstract
Bessel’s correction adjusts the denominator in the sample variance formula from $n$ to $n-1$ to ensure an unbiased estimator of the population variance. This paper provides rigorous algebraic derivations, geometric interpretations, and visualizations to reinforce the necessity of this correction. It further introduces the concept of Bariance, an alternative dispersion measure based on pairwise squared differences that avoids reliance on the arithmetic mean. Building on this, we address practical concerns raised in Rosenthal’s article [8], which advocates for $n$-based estimates from a mean squared error (MSE) perspective—particularly in pedagogical contexts and specific applied settings. Finally, the empirical component of this work, based on simulation studies, demonstrates that estimating the population variance via an algebraically optimized Bariance approach can yield a computational advantage. Specifically, the unbiased Bariance estimator can be computed in linear time, resulting in shorter run-times while preserving statistical validity.
JEL Codes: C10, C80
Keywords: Unbiased sample variance, Runtime-optimized linear unbiased sample variance estimators
“This could be the end of it all,
released from the pain.
So self-indulgent and so insincere,
I’ll never bow to your lies again.”
— Megadeth, *Tornado of Souls*
The following notation is used throughout the paper:
$X_{1},X_{2},\dots,X_{n}\in\mathbb{R}$
Independent and identically distributed sample from a population.
$\mu=\operatorname{\mathbb{E}}[X_{i}]$ Population mean.
$\sigma^{2}=\operatorname{\mathrm{Var}}(X_{i})$ Population variance.
$\bar{X}=\frac{1}{n}\sum_{i=1}^{n}X_{i}$ Sample mean.
$S^{2}=\frac{1}{n}\sum_{i=1}^{n}(X_{i}-\bar{X})^{2}$ Biased sample variance estimator (denominator $n$).
$\hat{S}^{2}=\frac{1}{n-1}\sum_{i=1}^{n}(X_{i}-\bar{X})^{2}$ Unbiased sample variance estimator (Bessel-corrected).
$\text{{Bariance}}=\frac{1}{n(n-1)}\sum_{i\neq j}(X_{i}-X_{j})^{2}$ Pairwise variance estimator using all ordered pairs (each unordered pair counted twice).
$\text{{Bariance}}_{\text{opt}}=\frac{2n}{n(n-1)}\sum_{i=1}^{n}X_{i}^{2}-\frac{2}{n(n-1)}\left(\sum_{i=1}^{n}X_{i}\right)^{2}$ Optimized scalar formula for Bariance.
$\sum_{i<j}(X_{i}-X_{j})^{2}$ Sum over all unordered pairwise squared differences (each pair counted once).
Note: $\sum_{i\neq j}(X_{i}-X_{j})^{2}=2\sum_{i<j}(X_{i}-X_{j})^{2}$.
$\text{Bias}(\hat{\theta})=\operatorname{\mathbb{E}}[\hat{\theta}]-\theta$ Bias of an estimator $\hat{\theta}$.
$\text{Var}(\hat{\theta})=\operatorname{\mathbb{E}}[(\hat{\theta}-\operatorname{\mathbb{E}}[\hat{\theta}])^{2}]$ Variance of an estimator.
$\text{MSE}(\hat{\theta})=\text{Var}(\hat{\theta})+\text{Bias}^{2}(\hat{\theta})$ Mean squared error of an estimator.
$\vec{X}\in\mathbb{R}^{n}$ Sample vector in Euclidean space.
$\vec{1}\in\mathbb{R}^{n}$ All-ones vector.
$\vec{r}=\vec{X}-\vec{\mu}$ Residual vector after projection onto the mean.
$\vec{r}\in\vec{1}^{\perp}$ Residual lies in the orthogonal complement of the mean vector.
$\Gamma(k,\theta)$ Gamma distribution with shape $k$ and scale $\theta$.
$\mathcal{N}(\mu,\sigma^{2})$ Normal distribution with mean $\mu$ and variance $\sigma^{2}$.
$\sum X_{i}$ Sum of sample values.
$\sum X_{i}^{2}$ Sum of squared sample values.
Note: The term Bariance emphasizes variance computed from pairwise squared differences instead of deviations from the mean.
1 Introduction and Motivation
Variance estimation is a foundational task in statistics and econometrics, with the sample variance being the default estimator in most applications. The unbiased version, corrected by Bessel’s factor (dividing by $n-1$ rather than $n$), compensates for the loss of one degree of freedom due to pre-estimating the population mean. This correction is not just a simple algebraic trick—it admits deep geometric interpretations via orthogonal projections in $\mathbb{R}^{n}$ and can be derived rigorously from them.
Despite its theoretical appeal, the unbiased estimator is not always the most optimal in practice. In small samples especially, its higher variance may lead to suboptimal inference. This has led researchers to consider shrunken estimators that intentionally trade off a small amount of bias for a significant reduction in variance, thereby minimizing mean squared error (MSE). For example, empirical Bayes methods shrink sample variances toward a global prior, stabilizing estimation across thousands of features in genomic studies [9]. Similar techniques based on James–Stein shrinkage have been explored for variance estimation in high-dimensional settings [4].
Beyond the univariate case, shrinkage ideas are especially powerful in multivariate settings. In particular, shrinkage estimators for covariance matrices—such as the Ledoit–Wolf estimator [6]—have gained popularity in fields like econometrics and finance, particularly in the field of asset pricing. These estimators enhance the stability of sample covariance matrices by shrinking them toward structured targets (e.g., the identity matrix), significantly improving conditioning in high-dimensional models, which are known to perform poorly [5]. This has practical relevance in the construction of variance-covariance matrices for portfolio optimization, factor models, and robust standard error estimation in large-scale regression analysis for econometric applications.
In this broader context, this paper revisits classical variance estimation and introduces a novel perspective via an alternative measure of sample dispersion based on the average squared differences between all unordered pairs in a sample. We formally define this estimator as the Bariance, a term that reflects its construction from pairwise distances rather than deviations from a mean. It can be shown that for mean-centered data, the Bariance equals exactly twice the unbiased sample variance. Moreover, a linear-time optimized formulation of the Bariance can be derived using simple algebraic properties that avoids quadratic pairwise computation, making it both theoretically elegant and computationally efficient.
Although the pairwise difference approach has roots in classical statistics—such as U-statistics [7], dissimilarity-based dispersion measures, and even the Gini coefficient [2]—the contribution here is a novel, unbiased estimator that is computationally optimized for runtime efficiency. In this respect, Bariance bridges theoretical variance estimation with algorithmic efficiency, a consideration critical in big data contexts, real-time systems, and streaming analytics.
While computational efficiency is one of its key advantages, the Bariance measure may also prove valuable in applied scenarios where the concept of central tendency is unstable, ill-defined, or misleading. For example, in domains such as network analysis, genomics, ordinal survey research, or clustering, statistical dispersion is often better captured through relational or pairwise structures rather than deviations from a single global mean. In such contexts, the Bariance shares conceptual kinship with the Gini coefficient, which also operates on pairwise differences but in a distributional inequality framework. Unlike Gini, however, Bariance preserves unbiasedness for variance estimation under i.i.d. sampling and scales naturally in high-dimensional or streaming environments. These features make it particularly attractive for modern applications in unsupervised learning, robust statistics, and high-throughput data pipelines—where traditional variance measures may either fail or become computationally prohibitive.
Through an empirical simulation study, I demonstrate that this optimized unbiased sample variance estimator remains unbiased and improves runtime. The simulated empirical runtimes section includes confidence intervals, hardware specifications, seed initialization, and multiple replications, thereby addressing robustness, reproducibility, and statistical reliability. Furthermore, Appendix C replicates the result in a local hardware environment using an alternative high-level programming language. We then revisit the controversial idea—advocated by Rosenthal [8]—that dividing by $n$ (rather than $n-1$) may yield lower-MSE variance estimators in practice, especially when unbiasedness is not strictly required.
To sum up, the Bariance framework bridges computational efficiency with applied relevance, offering a theoretically grounded yet practically flexible alternative to traditional variance estimators. This paper thus aims to bridge classical econometric and statistical theory with modern considerations of efficiency, robustness, and computational scalability, while highlighting the often underestimated choices in estimator design or usage.
2 Definitions and Setup
Let $X_{1},X_{2},\dots,X_{n}\in\mathbb{R}$ be i.i.d. random variables with:
$\operatorname{\mathbb{E}}[X_{i}]=\mu,\qquad\operatorname{\mathrm{Var}}(X_{i})=\sigma^{2}$
Define the sample mean and biased/unbiased variance estimates:
$\bar{X}:=\frac{1}{n}\sum_{i=1}^{n}X_{i},\quad S^{2}:=\frac{1}{n}\sum_{i=1}^{n}(X_{i}-\bar{X})^{2},\quad\hat{S}^{2}:=\frac{1}{n-1}\sum_{i=1}^{n}(X_{i}-\bar{X})^{2}$
3 Derivation of Bias and Bessel’s Correction
An estimator $\hat{\theta}$ for a parameter $\theta$ is called unbiased if its expected value equals the true value:
$\operatorname{\mathbb{E}}[\hat{\theta}]=\theta$
The normal $n$-based sample variance with denominator $n$ is defined as:
$S^{2}:=\frac{1}{n}\sum_{i=1}^{n}(X_{i}-\bar{X})^{2}$
We aim to compute $\operatorname{\mathbb{E}}[S^{2}]$, the expected value of this estimator, to show that it is biased.
Expand the squared deviations:
$\sum_{i=1}^{n}(X_{i}-\bar{X})^{2}\equiv\sum_{i=1}^{n}X_{i}^{2}-n\bar{X}^{2}$
Thus:
$S^{2}=\frac{1}{n}\left(\sum_{i=1}^{n}X_{i}^{2}-n\bar{X}^{2}\right)=\frac{1}{n}\sum X_{i}^{2}-\bar{X}^{2}$
Then, take expectation of $S^{2}$. By linearity of expectation to each term:
$\operatorname{\mathbb{E}}[S^{2}]=\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathbb{E}}[X_{i}^{2}]-\operatorname{\mathbb{E}}[\bar{X}^{2}]$
Compute $\operatorname{\mathbb{E}}[X_{i}^{2}]$. Using the known identity:
$\operatorname{\mathbb{E}}[X_{i}^{2}]\equiv\operatorname{\mathrm{Var}}(X_{i})+(\operatorname{\mathbb{E}}[X_{i}])^{2}=\sigma^{2}+\mu^{2}$
So, the $n$ cancels out, eventually:
$\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathbb{E}}[X_{i}^{2}]=\frac{1}{n}\cdot n(\mu^{2}+\sigma^{2})=\mu^{2}+\sigma^{2}$
Compute $\operatorname{\mathbb{E}}[\bar{X}^{2}]$. Recall that:
$\bar{X}=\frac{1}{n}\sum_{i=1}^{n}X_{i}\quad\Rightarrow\quad\operatorname{\mathbb{E}}[\bar{X}]=\mu,\quad\operatorname{\mathrm{Var}}(\bar{X})=\frac{\sigma^{2}}{n}$
Thus:
$\operatorname{\mathbb{E}}[\bar{X}^{2}]=\operatorname{\mathrm{Var}}(\bar{X})+(\operatorname{\mathbb{E}}[\bar{X}])^{2}=\frac{\sigma^{2}}{n}+\mu^{2}$
Combining both terms now:
$\operatorname{\mathbb{E}}[S^{2}]=(\mu^{2}+\sigma^{2})-\left(\mu^{2}+\frac{\sigma^{2}}{n}\right)=\sigma^{2}-\frac{\sigma^{2}}{n}=\left(\frac{n-1}{n}\right)\sigma^{2}$
$\boxed{\operatorname{\mathbb{E}}[S^{2}]=\frac{n-1}{n}\sigma^{2}}$
This shows that the estimator $S^{2}$ is biased, underestimating the population variance $\sigma^{2}$, because the denominator is larger than the numerator.
Bessel’s Correction. To correct the bias, we define the unbiased sample variance as:
$\hat{S}^{2}:=\frac{1}{n-1}\sum_{i=1}^{n}(X_{i}-\bar{X})^{2}\Rightarrow\operatorname{\mathbb{E}}[\hat{S}^{2}]=\sigma^{2}$
[…]
This is known as Bessel’s correction — using $n-1$ instead of $n$ in the denominator compensates for the loss of one degree of freedom from estimating the mean $\mu$ with $\bar{X}$.
4 Geometric Interpretation of Estimated Variance and $n-1$ Degrees of Freedom
Orthogonal Decomposition
Let $\vec{X}:=\begin{bmatrix}X_{1}&\dots&X_{n}\end{bmatrix}^{\top}\in\mathbb{R}^{n}$. Define the mean
$\bar{X}=\frac{1}{n}\sum_{i=1}^{n}X_{i},\quad\vec{\mu}=\bar{X}\cdot\vec{1}.$
Then
$\vec{X}=\vec{\mu}+\vec{r},\quad\vec{r}:=\vec{X}-\vec{\mu},\quad\vec{r}\in\vec{1}^{\perp}.$
Indeed, $\vec{1}^{\top}\vec{r}=\sum_{i=1}^{n}(X_{i}-\bar{X})=0$.
[…]
Figure 1: Orthogonal decomposition of $\vec{X}$
Dimension and Degrees of Freedom
$\vec{X}$ $\in\mathbb{R}^{n}$
$\vec{\mu}$ $\in\operatorname{span}(\vec{1}),\quad\dim=1$
$\vec{r}$ $\in\vec{1}^{\perp},\quad\dim=n-1$
[…]
Unbiased Sample Variance
$s^{2}=\frac{1}{n-1}\|\vec{r}\|^{2}=\frac{1}{n-1}\sum_{i=1}^{n}(X_{i}-\bar{X})^{2}$
Common Application using Orthogonal Decomposition for Dimensionality reduction: Principal Component Analysis (PCA)
Let $\mathbf{X}\in\mathbb{R}^{n\times p}$ be a data matrix with rows as observations and columns as variables. Let $\tilde{X}:=X-\vec{1}\bar{x}^{\top}$, where $\bar{x}=\frac{1}{n}X^{\top}\vec{1}\in\mathbb{R}^{p}$.
PCA seeks orthonormal vectors $\vec{v}_{i}\in\mathbb{R}^{p}$ satisfying:
$\tilde{X}^{\top}\tilde{X}\vec{v}_{i}=\lambda_{i}\vec{v}_{i},\quad i=1,\dots,p.$
Then $\vec{z}_{i}:=\tilde{X}\vec{v}_{i}$ are commonly known as $i$th principal components ($PC_{i}$), with variances $\operatorname{Var}(\vec{z}_{i})=\lambda_{i}/(n-1)$.
5 Introducing the Bariance and an optimized linear Estimator
We define the Bariance of a sample $\{X_{1},X_{2},\dots,X_{n}\}$ as the average squared difference over all unordered pairs:
$\text{{Bariance}}:=\frac{1}{n(n-1)}\sum_{i\neq j}(X_{i}-X_{j})^{2}$
The term Bariance is simply chosen to emphasize the estimator’s foundation on pairwise between-sample variance rather than deviations from the mean, highlighting its construction from pairwise squared differences. This can also be interpreted as the average squared length of all edges in a complete graph on the sample points.
We begin by expanding the inner squared difference:
$(X_{i}-X_{j})^{2}=X_{i}^{2}-2X_{i}X_{j}+X_{j}^{2}$
Summing over all distinct $i\neq j$:
$\sum_{i\neq j}(X_{i}-X_{j})^{2}=\sum_{i\neq j}(X_{i}^{2}+X_{j}^{2}-2X_{i}X_{j})$
We split this into three terms:
$=\sum_{i\neq j}X_{i}^{2}+\sum_{i\neq j}X_{j}^{2}-2\sum_{i\neq j}X_{i}X_{j}$
Note the following observations: - For fixed $i$, there are $n-1$ values of $j\neq i$, so:
$\sum_{i\neq j}X_{i}^{2}=(n-1)\sum_{i=1}^{n}X_{i}^{2}$
Similarly, $\sum_{i\neq j}X_{j}^{2}=(n-1)\sum_{j=1}^{n}X_{j}^{2}$
So the first two terms become:
$\sum_{i\neq j}X_{i}^{2}+\sum_{i\neq j}X_{j}^{2}=2(n-1)\sum_{i=1}^{n}X_{i}^{2}$
Now consider the double sum:
$\sum_{i\neq j}X_{i}X_{j}=\left(\sum_{i=1}^{n}\sum_{j=1}^{n}X_{i}X_{j}\right)-\sum_{i=1}^{n}X_{i}^{2}=\left(\sum X_{i}\right)^{2}-\sum X_{i}^{2}$
Combine:
$\sum_{i\neq j}(X_{i}-X_{j})^{2}=2(n-1)\sum X_{i}^{2}-2\left(\left(\sum X_{i}\right)^{2}-\sum X_{i}^{2}\right)$
$=2(n-1)\sum X_{i}^{2}-2\left(\sum X_{i}\right)^{2}+2\sum X_{i}^{2}=2n\sum X_{i}^{2}-2\left(\sum X_{i}\right)^{2}$
Substitute back into the Bariance formula. Now divide by $n(n-1)$:
$\text{{Bariance}}=\frac{1}{n(n-1)}\sum_{i\neq j}(X_{i}-X_{j})^{2}\equiv\frac{2n}{n(n-1)}\sum X_{i}^{2}-\frac{2}{n(n-1)}\left(\sum X_{i}\right)^{2}.$
For empirical verification of this algebraic identity see Appendix A.
$\boxed{\text{{Bariance}}_{\text{opt}}:=\frac{2n}{n(n-1)}\sum X_{i}^{2}-\frac{2}{n(n-1)}\left(\sum X_{i}\right)^{2}.}$
5.0.1 In the Case Of Mean-centered data
If the data is centered, i.e., $\sum X_{i}=0$, then:
$\text{{Bariance}}=\frac{2n}{n(n-1)}\sum X_{i}^{2}=\frac{2}{n-1}\sum X_{i}^{2}.$
We now relate this to the unbiased sample variance estimator:
$\hat{S}^{2}=\frac{1}{n-1}\sum_{i=1}^{n}(X_{i}-\bar{X})^{2}=\frac{1}{n-1}\sum X_{i}^{2}\quad\text{(since }\bar{X}=0\text{)}.$
Therefore, the following equality holds for the defined “Bariance“:
$\boxed{\text{{Bariance}}=2\cdot\hat{S}^{2}.}$
This result shows that Bariance represents twice the unbiased sample variance when the sample is mean-centered. It provides an elegant pairwise perspective on variance: instead of summing squared deviations from a central value, we sum squared differences between all pairs and average, regardless of the reference point within the sample.
5.1 Properties of the Bariance and Estimator Comparison
Let $\theta:=\sigma^{2}=1$. Then:
$\operatorname{\mathbb{E}}[\hat{S}^{2}]$ $=\theta,$
$\operatorname{\mathbb{E}}[\text{{Bariance}}]$ $=2\theta,$
$\text{Bias}(\hat{S}^{2})$ $=0,$
[…]
$\operatorname{\mathrm{Var}}(\hat{S}^{2})$ $=\frac{2\theta^{2}}{n-1},$
$\operatorname{\mathrm{Var}}(\text{{Bariance}})$ $=4\cdot\operatorname{\mathrm{Var}}(\hat{S}^{2}),$
$\text{MSE}(\hat{S}^{2})$ $=\operatorname{\mathrm{Var}}(\hat{S}^{2}),$
[…]
Subsequent numerical verification of theoretical relationships with $n=100$ and $\tau=1000$
Numerically (for $n=100$, $\mathcal{N}(0,1)$, $\tau=1000$):
Estimator Point Estimate Bias Variance MSE
Unbiased sample variance estimator 1.00091 0.00091 0.02156 0.02151
Bariance 2.00181 1.00181 0.08625 1.08968
$\operatorname{\mathrm{Var}}(\hat{S}^{2})=0.01988,\quad\operatorname{\mathrm{Var}}(2\cdot\hat{S}^{2})=0.07951,$
$\text{MSE}(\hat{S}^{2})=0.02151,\quad\text{MSE}(2\cdot\hat{S}^{2})=1.0832,$
$4\cdot\operatorname{\mathrm{Var}}(\hat{S}^{2})=0.07951\quad\Rightarrow\quad\text{Var identity approximately holds},$
$4\cdot\operatorname{\mathrm{Var}}(\hat{S}^{2})+\theta^{2}=0.07951+1=1.07951\quad\Rightarrow\quad\text{MSE identity approximately holds}.$
[…]
Summary:
  • $\hat{S}^{2}$: unbiased, lower MSE, standard estimator.
  • Bariance: biased (by $+\theta$), higher MSE due to both variance inflation and squared bias.
  • Useful relation: $\text{{Bariance}}=2\cdot\hat{S}^{2}$. (Because: $\theta$ = ($\theta$ + $\theta$)/2)
5.2 A Graph-Theoretic View of Bariance
[…]
Figure 2: Complete graph over sample points. Each edge represents a pairwise squared difference that contributes to the Bariance.
Each dashed edge on the graph corresponds to a pair $(X_{i},X_{j})$, and the squared difference $(X_{i}-X_{j})^{2}$ can be viewed as an edge weight. Since the graph is fully connected (a complete graph $G_{n}$), there are $\binom{n}{2}$ such edges.
From a graph-theoretic standpoint, Bariance is the average squared edge length of the complete weighted graph over the sample. In this context:
- Nodes = observations $X_{i}$
- Edges = pairwise differences
- Weights = squared differences $(X_{i}-X_{j})^{2}$
This perspective connects naturally to distance-based dispersion measures in statistical graph theory, including energy statistics and certain U-statistics. It also provides an intuitive, coordinate-free alternative to the variance’s reliance on a central location (i.e., the mean).
5.3 Deviation from Mean (Variance) vs. Pairwise Differences (Bariance Components)
[…]
Figure 3: Blue: variance (mean-deviation, $n-1$ degrees of freedom adjustment). Green: pairwise distance = a Bariance component.
5.4 The Pairwise Difference Grid
0.04.016.036.064.04.00.04.016.036.016.04.00.04.016.036.016.04.00.04.064.036.016.04.00.0224466881010
Figure 4: Symmetric matrix of squared pairwise differences $(X_{i}-X_{j})^{2}$ for $X=\{2,4,6,8,10\}$. Diagonal elements are zero (self-distances).
This matrix visualizes all pairwise squared differences between elements of $X$. Its key properties:
- Symmetry: $(X_{i}-X_{j})^{2}=(X_{j}-X_{i})^{2}$, so the grid is symmetric across the main diagonal.
- Diagonal: Always zero, since $(X_{i}-X_{i})^{2}=0$.
- Off-diagonal structure: Each non-zero entry contributes to the total in the Bariance calculation.
This grid-based view reinforces how the redundancy in the pairwise differences allows us to simplify the computation algebraically. Instead of evaluating all $n(n-1)$ pairs, we can summarize the matrix using aggregate row/column sums, which enables the linear-time algebraically optimized derived formula for the Bariance:
$\text{{Bariance}}_{\text{opt}}=\frac{2n}{n(n-1)}\sum X_{i}^{2}-\frac{2}{n(n-1)}\left(\sum X_{i}\right)^{2}.$
Thus, symmetry is not just a visual feature—it underpins the algebraic transformation that reduces the quadratic computational complexity to linear.
5.5 Numerical Verification of Bariance-Estimator properties using Gamma-distributed Data
Let $X\sim\Gamma(k=2,\theta=2)\Rightarrow\operatorname{\mathbb{E}}[X]=4,\quad\operatorname{\mathrm{Var}}(X)=8$, Using $n=100$, $\tau=1000$ we obtain:
Summary Table of Empirical Bariance Point Estimator Performance
Estimator Mean Bias Variance MSE
Unbiased Sample Variance Estimator 8.00087 0.00087 2.92574 2.92281
Bariance 16.00174 8.00174 11.70295 75.71907
Numerical Verification of Theoretical Relationships
Let $\sigma^{2}:=8\Rightarrow\sigma^{4}=64$:
$\operatorname{\mathrm{Var}}(\hat{S}^{2})=2.92574,\quad\operatorname{\mathrm{Var}}(\text{{Bariance}})=11.70295,$
[…]
We numerically verify:
$\operatorname{\mathrm{Var}}(\text{{Bariance}})\approx 4\cdot\operatorname{\mathrm{Var}}(\hat{S}^{2})\quad\Rightarrow\quad 11.70295\approx 4\cdot 2.92574=11.70296,$
$\text{MSE}(\text{{Bariance}})\approx 4\cdot\operatorname{\mathrm{Var}}(\hat{S}^{2})+\sigma^{4}\quad\Rightarrow\quad 75.71907\approx 4\cdot 2.92574+64=75.70295,$
[…]
6 Discussion: Should We Just Divide by $n$?
Rosenthal [8] argues that using $n$ instead of $n-1$ may lead to a smaller mean squared error (MSE) — especially when teaching or in practical settings.
He shows that while dividing by $n-1$ yields an unbiased estimator, this may come at the cost of increased variance. In some cases, a biased but lower-MSE estimator using $n$ is preferable:
“…a smaller, shrunken, biased estimator actually reduces the MSE…” — [8]
This introduces another viewpoint: unbiasedness isn’t always the ultimate goal — minimizing error in practice often is.
From a theoretical perspective, unbiasedness ensures that the expected value of the estimator exactly matches the true population variance. However, unbiasedness alone does not guarantee minimal estimation error in finite samples. In fact, particularly when sample sizes are small, the variance of the unbiased estimator can be relatively large, which may lead to unstable or inefficient estimates. Allowing a small bias can reduce this variance enough to produce a lower overall mean squared error, which combines both bias and variance in a single measure of estimator quality [1, 10].
Practically, this trade-off becomes important in many applied contexts, such as when the variance estimate is an intermediate quantity used for further modeling or prediction. Here, reducing total estimation error takes precedence over strict unbiasedness. Additionally, computational simplicity and pedagogical clarity sometimes favor the $n$-denominator estimator, which is easier to understand and implement, especially in introductory settings.
To illustrate, consider the case of $n=5$ observations drawn from a population with true variance $\sigma^{2}=10$. The biased estimator, which divides by $n=5$, underestimates the variance as $8$, while the unbiased estimator, dividing by $n-1=4$, correctly yields $10$. However, when mean squared error is calculated, which accounts for both bias and variance, the biased estimator can have smaller total error due to its reduced variance. Rosenthal explicitly notes that the variance of the biased estimator is often less than that of the unbiased one, which explains this result.
This nuanced understanding clarifies why and when the classical insistence on unbiasedness may be relaxed in favor of better finite-sample performance and practical utility. Furthermore, more generalized estimators with a denominator parameter $a>0$, defined as
$\hat{\sigma}^{2}_{a}:=\frac{1}{a}\sum_{i=1}^{n}(X_{i}-\bar{X})^{2},$
can be analyzed for optimality in terms of mean squared error. As derived in Appendix B, the MSE-minimizing denominator is approximately $a^{*}\approx n+1$, suggesting that neither the classical unbiased divisor $n-1$ nor the biased $n$ divisor are necessarily optimal from an MSE standpoint.
In sum, relaxing unbiasedness for variance estimation is a principled and context-dependent choice motivated by the bias-variance trade-off, sample size considerations, and the ultimate goals of estimation. This perspective complements classical theory and better reflects the realities of applied statistical practice.
7 A Simulation Study: Bias2, Variance, and MSE Across Denominator Values
We consider the family of estimators for the population variance $\sigma^{2}$:
$\hat{\sigma}^{2}_{a}:=\frac{1}{a}\sum_{i=1}^{n}(X_{i}-\bar{X})^{2}\quad\text{for varying }a>0$
The simulation is carried out with the following parameters:
  • Sample size: $n=5$
  • True variance: $\sigma^{2}=10$
  • Distribution: $X_{i}\sim\mathcal{N}(0,\sigma^{2})$
  • Number of simulations: $100{,}000$
For each value of $a\in[3.5,8.5]$ (in increments of 0.5), we compute the following empirically:
$\text{Bias}(\hat{\sigma}^{2}_{a})$ $=\mathbb{E}[\hat{\sigma}^{2}_{a}]-\sigma^{2}$
$\text{Bias}^{2}$ $=\left(\mathbb{E}[\hat{\sigma}^{2}_{a}]-\sigma^{2}\right)^{2}$
Variance $=\text{Var}[\hat{\sigma}^{2}_{a}]$
MSE $=\text{Bias}^{2}+\text{Variance}$
Empirical Results
Table 1: Empirical Bias2, Variance, and MSE with 95% bootstrapped confidence intervals (200 resamples, seed=42). Bold rows indicate $a=n-1$, $n$, and $n+1$. Hardware: 1GB RAM, Python 3.11. For a theoretical derivation of the MSE-minimized variance estimator, see Appendix B.
$a$ Bias2 [CI] Variance [CI] MSE [CI]
3.5 1.98 [1.97, 2.03] 61.99 [61.85, 62.19] 63.96 [63.83, 64.21]
4.0 0.00 [0.00, 0.01] 47.46 [47.26, 47.51] 47.46 [47.26, 47.52]
4.5 1.27 [1.26, 1.30] 37.50 [37.34, 37.57] 38.77 [38.63, 38.84]
5.0 4.06 [4.02, 4.08] 30.37 [30.29, 30.47] 34.44 [34.36, 34.50]
5.5 7.52 [7.48, 7.54] 25.10 [25.07, 25.20] 32.62 [32.59, 32.70]
6.0 11.20 [11.16, 11.24] 21.09 [21.05, 21.17] 32.29 [32.27, 32.36]
6.5 14.89 [14.81, 14.90] 17.97 [17.92, 18.03] 32.86 [32.79, 32.87]
7.0 18.46 [18.46, 18.56] 15.50 [15.42, 15.51] 33.96 [33.94, 34.01]
7.5 21.88 [21.85, 21.94] 13.50 [13.48, 13.55] 35.37 [35.37, 35.44]
8.0 25.10 [25.06, 25.17] 11.86 [11.82, 11.89] 36.96 [36.93, 37.01]
8.5 28.13 [28.09, 28.20] 10.51 [10.45, 10.52] 38.64 [38.59, 38.67]
33333.53.53.53.544444.54.54.54.555555.55.55.55.566666.56.56.56.577777.57.57.57.588888.58.58.58.59999005555101010101515151520202020252525253030303035353535404040404545454550505050Denominator $a$ValueMSE, Bias2, and Variance of Variance EstimatorMSEBias2VarianceDivide by $n{-}1$Divide by $n$Divide by $n{+}1$
Figure 5: Empirical MSE, Bias2, and Variance of the sample variance estimator for $a\in[3.5,8.5]$ and $n=5$ using 10,000 simulations and 200 bootstraps. Minimum MSE occurs between $a=5.5$ and $a=6.5$. For a theoretical derivation of the MSE-minimized variance estimator, see Appendix B.
8 On Computational Complexity of Variance Bariance Estimators and Optimization
Let $X:=\{X_{1},X_{2},\dots,X_{n}\}\subset\mathbb{R}$ be a sample of size $n$.
Table 2: Computational complexity of variance and Bariance estimators with explanation
Estimator Operations Complexity
Biased Variance
$S^{2}=\frac{1}{n}\sum(X_{i}-\bar{X})^{2}$ 1 pass to compute mean $\bar{X}$ 1 pass to compute squared deviations Total: 2 linear scans For $n=5$: 5 additions, 5 subtractions, 5 squarings $\mathcal{O}(n)$
Unbiased Variance
$\hat{S}^{2}=\frac{1}{n-1}\sum(X_{i}-\bar{X})^{2}$ Same steps as biased estimator; only the divisor differs.
No added computation.
$\mathcal{O}(n)$
Bariance (Naïve)
$\frac{1}{n(n-1)}\sum_{i\neq j}(X_{i}-X_{j})^{2}$ All $n(n-1)$ ordered pairs evaluated Each requires subtraction + squaring For $n=5$: $5\times 4=20$ pairs Cost grows quadratically with sample size $\mathcal{O}(n^{2})$
Bariance (Optimized)
$\frac{2n}{n(n-1)}\sum X_{i}^{2}-\frac{2}{n(n-1)}\left(\sum X_{i}\right)^{2}$ Uses 2 scalar sums: $\sum X_{i}$, $\sum X_{i}^{2}$ Each computed in 1 pass For $n=5$: 5 additions, 5 squarings $\mathcal{O}(n)$
8.1 Computational Complexity Comparison with Numerical Illustration
We compare the computational cost of the biased variance, unbiased variance, and Bariance estimators using both theoretical analysis and a numerical example for $n=5$.
Example: $X=\{1,3,5,7,9\}$
Mean:
$\bar{X}=\frac{1+3+5+7+9}{5}=\frac{25}{5}=5$
Biased Variance:
$S^{2}=\frac{1}{5}\sum(X_{i}-\bar{X})^{2}=\frac{1}{5}[(1-5)^{2}+(3-5)^{2}+(5-5)^{2}+(7-5)^{2}+(9-5)^{2}]=\frac{1}{5}[16+4+0+4+16]=\frac{40}{5}=8$
Unbiased Variance:
$\hat{S}^{2}=\frac{1}{4}\sum(X_{i}-\bar{X})^{2}=\frac{40}{4}=10$
Naïve Bariance:
$\sum_{i<j}(X_{i}-X_{j})^{2}=(3-1)^{2}+(5-1)^{2}+(7-1)^{2}+(9-1)^{2}+(5-3)^{2}+(7-3)^{2}+(9-3)^{2}+(7-5)^{2}+(9-5)^{2}+(9-7)^{2}$
$=4+16+36+64+4+16+36+4+16+4=200$
$\text{{Bariance}}=\frac{2\cdot 200}{5\cdot 4}=\frac{400}{20}=20$
Optimized Bariance:
$\sum X_{i}=1+3+5+7+9=25,\quad\sum X_{i}^{2}=1^{2}+3^{2}+5^{2}+7^{2}+9^{2}=165$
$\text{{Bariance}}=\frac{2n}{n(n-1)}\sum X_{i}^{2}-\frac{2}{n(n-1)}\left(\sum X_{i}\right)^{2}=\frac{2\cdot 5}{20}\cdot 165-\frac{2}{20}\cdot 625=\frac{1650}{20}-\frac{1250}{20}=82.5-62.5=20$
Thus, all estimators yield consistent results, confirming their correctness. However, their computational complexity differs:
  • Biased/Unbiased Variance: $\mathcal{O}(n)$
  • Naïve Bariance: $\mathcal{O}(n^{2})$
  • Optimized Bariance: $\mathcal{O}(n)$
The optimized Bariance offers the same result as the naïve form but with significantly reduced computational cost, making it efficient for large-scale applications.
9 On Empirical Runtime of Bariance Estimators
To evaluate the practical performance of variance and Bariance estimators, we conducted an empirical benchmark based on simulated data. The goal was to measure actual computation time across increasing sample sizes for the four as above defined estimators.
9.1 Execution Environment
All Python code was executed in a virtualized Python 3.11.8 environment on a Linux system (kernel version 4.4.0) with x86_64 architecture. The processor was identified as unknown, featuring 32 logical cores and 32 physical cores. The system reported a BogoMIPS value of 2593.91.
Available memory was limited to 1.07 GB of RAM, with no swap configured. Supported CPU instruction sets included AVX, AVX2, AVX-512, and FMA, as listed in /proc/cpuinfo.
All timing measurements were performed using time.perf_counter to capture high-resolution wall-clock time in single-threaded execution mode. Each estimator was evaluated across $\tau$ = 20 independent trials per sample size.
9.2 Normal-Distributed Data
  • Number of simulations per sample size: 1000
  • Sample sizes tested: $n\in\{10,20,\dots,100\}$
  • Distribution: $X_{i}\sim\mathcal{N}(0,1)$
  • Timing measurement: Wall-clock time per estimator (summed over 1000 replications)
All implementations were naïvely vectorized using broadcasting or looped to mimic real computational effort and make the comparison fair between estimator types.
Table 3: Empirical runtime (in seconds) for 1,000 simulations per estimator across different sample sizes ($n$). Time measured in wall-clock seconds. Bold values indicate the fastest method for each row.
$n$ Biased Variance Unbiased Variance Bariance (Naïve) Bariance (Optimized)
10 0.0131 0.0142 0.0601 0.0119
20 0.0208 0.0143 0.2191 0.0092
30 0.0115 0.0115 0.4872 0.0091
40 0.0121 0.0123 0.8767 0.0104
50 0.0134 0.0132 1.5155 0.0092
60 0.0124 0.0122 2.1050 0.0090
70 0.0186 0.0176 2.7712 0.0087
80 0.0126 0.0205 3.6592 0.0155
90 0.0139 0.0135 5.0322 0.0095
100 0.0127 0.0125 5.6617 0.0098
9.3 Gamma-Distributed Data
To examine runtime behavior under non-Gaussian conditions, we conducted a second simulation study using data generated from a Gamma distribution. The $\Gamma$-distribution is positively skewed, making it a useful alternative to test estimator performance beyond the symmetric $\mathcal{N}$ case.
Parameters of the Gamma-Based Simulation
  • Number of simulations per sample size: 500
  • Sample sizes tested: $n\in\{100,200,300,400,500\}$
  • Distribution: $X_{i}\sim\Gamma(2,2)$
  • Timing measurement: Wall-clock time per estimator (summed over 500 replications)
Table 4: Empirical runtime (in seconds) for 500 simulations per estimator using Gamma-distributed data. Time measured in wall-clock seconds. Bold values highlight the fastest method at each sample size $n$.
$n$ Biased Variance Unbiased Variance Bariance (Naïve) Bariance (Optimized)
100 0.0073 0.0105 0.0149 0.0065
200 0.0083 0.0101 0.0430 0.0084
300 0.0080 0.0102 0.1075 0.0073
400 0.0077 0.0101 0.1937 0.0074
500 0.0128 0.0164 0.3266 0.0095
9.4 Highly Dispersed Gamma-Distributed Data
To further assess runtime robustness under high skew and dispersion, we generated data from a $\Gamma$ distribution with increased variance. This setup simulates conditions with greater variability, which are common in skewed real-world datasets.
Parameters of the Highly Dispersed Gamma-Based Simulation
  • Number of simulations per sample size: 1000
  • Sample sizes tested: $n\in\{50,100,150,200,250\}$
  • Distribution: $X_{i}\sim\Gamma(1.5,4.0)$
  • Timing measurement: Wall-clock time per estimator (summed over 1000 replications)
Table 5: Empirical runtime (in seconds) for 1,000 simulations per estimator using highly dispersed Gamma-distributed data. Time measured in wall-clock seconds. Bold values indicate the fastest method for each sample size $n$.
$n$ Biased Variance Unbiased Variance Bariance (Naïve) Bariance (Optimized)
50 0.0134 0.0171 0.0173 0.0141
100 0.0132 0.0171 0.0284 0.0121
150 0.0139 0.0184 0.0507 0.0128
200 0.0156 0.0183 0.0831 0.0127
250 0.0161 0.0179 0.1284 0.0129
9.5 Robustness: Statistical Analysis of Empirical Runtime
For Robustness testing in an alternative execution environment see Appendix C and the replication package  [^1] https://github.com/felix-reichel/BarianceVariance_Reproduction_Repo_Robustness_Supplements.
9.5.1 Execution Environment
Available memory was limited to 1.07 GB of RAM, with no swap configured. Supported CPU instruction sets included AVX, AVX2, AVX-512, and FMA, as listed in /proc/cpuinfo.
All timing measurements were performed using time.perf_counter to capture high-resolution wall-clock time in single-threaded execution mode. Each estimator was evaluated across $\tau$ = 20 independent trials per sample size.
9.5.2 Simulation Setup
All estimators were tested on synthetic data generated from a standard normal distribution $\mathcal{N}(0,1)$. The following estimators were implemented:
  • Biased and Unbiased Variance (Looped and Vectorized)
  • Optimized Bariance (Looped and Vectorized)
Sample sizes ranged from 100 to 4800 in even steps for a total of 48 tests, with an additional focused study at 16 evenly spaced sizes for detailed visualization. See Figure 7.
9.5.3 Main Estimator Comparison
Figure 6 shows kernel density estimates of runtime differences between the Unbiased Vectorized and Optimized Bariance Vectorized estimators across 16 sample sizes. The densities summarize performance variability over multiple runs.
Figure 6: Kernel density estimates of runtime differences between Unbiased Sample Variance Vectorized and Optimized Bariance Vectorized estimators. Positive values indicate slower performance of the unbiased estimator.
9.5.4 Kernel Density Comparison: 3x16 Grid
Figure 7 presents a 3x16 grid of kernel density estimates for runtime differences between the same two estimators across 48 sample sizes. Each panel shows the distribution for a distinct sample size $n$.
Figure 7: Kernel density estimates of runtime differences for Unbiased Sample Variance Vectorized and Optimized Bariance Vectorized estimators across 48 sample sizes.
9.5.5 All Estimators: Comparative Runtime Analysis
Figure 8 displays kernel density estimates of runtime for all estimators over 16 sample sizes. The two primary estimators are emphasized in color, while others appear in grayscale for contrast.
Figure 8: Kernel density estimates of runtime for all estimators across 16 sample sizes. Unbiased Sample Variance Vectorized and Optimized Bariance Vectorized are highlighted; grayscale lines represent additional looped and vectorized variants.
10 Conclusion
Bessel’s correction is a foundational concept that ensures unbiased estimates of variance. We explored its necessity through algebraic, geometric, and pairwise differences reasoning (now formalized as the Bariance construct), building both intuition and understanding. Additionally, we considered a pedagogical and practical perspective, such as Rosenthal’s $MSE$-based view for estimating variance [8].
Although the unbiased estimator is mathematically correct in expectation, the biased version can sometimes be more intuitive and, in certain contexts, statistically preferable across various sampling distributions. This aligns with insights from modern treatments of mathematical statistics, which often emphasize the trade-off between bias and variance in estimator performance [1, 10]. Furthermore, empirical results revealed a faster runtime in our simulation example using the average pairwise differences definition as an unbiased variance estimator—referred to as the Bariance estimator—particularly when employing the algebraically optimized formula using scalar sums.
To sum up, the main finding—the run-time optimized estimator for the Bariance formula—was a coincidental yet significant observation: that is, the unbiased estimator can be computed in linear time and statistically outperforms the conventional unbiased sample variance estimator in all tested empirical runtime performance scenarios. Naturally, many other estimators exist for sample variance, including those designed to trade off bias for computational gains. A complexity theorist or mathematician could potentially derive theoretical bounds on the time complexity of such estimators.
Beyond its computational efficiency, however, the Bariance measure may also offer substantive benefits in applied settings where deviation from a central mean is either unstable, undefined, or conceptually inappropriate. In fields such as genomics, network analysis, robust statistics, or ordinal survey research, dispersion may be more meaningfully characterized by average pairwise differences rather than deviations from a global average. Moreover, distance-based methods like clustering, energy statistics, and nonparametric ANOVA can all benefit from the geometric and symmetry-preserving properties of Bariance, particularly in high-dimensional or irregularly structured data where the mean offers little interpretive value. These contexts highlight how the pairwise construction of Bariance is not only computationally attractive but also methodologically appropriate.
Thus, the optimized Bariance formula stands as a viable alternative with promising practical implications for real-time multivariate big data applications, including forecasting (especially with shrunken variance-covariance estimators), computational biology, chemistry, finance, and big-data streaming applications (such as online learning) where unbiased and scalable variance estimation is essential.
Appendix A Proof of Equivalence: Naïve vs Optimized Bariance Estimators
Proof.
To verify the theoretical equivalence between the naïve and optimized formulations of the Bariance estimator, we conducted a simulation study using the exact formulas defined in Table 2. The data were drawn from a highly dispersed $\Gamma$- distribution.
Estimator Formulas
  • Naïve Bariance:
[…]
  • Optimized Bariance:
    $\text{{Bariance}}_{\text{opt}}=\frac{2n}{n(n-1)}\sum X_{i}^{2}-\frac{2}{n(n-1)}\left(\sum X_{i}\right)^{2}$
Simulation Parameters
  • Distribution: $\Gamma$($1.5$, $4.0$)
  • Sample sizes: $n\in\{50,100,150,200,250\}$
  • Number of simulations per $n$: 1000
  • Language: Python (NumPy)
  • Precision check: numpy.allclose with $\texttt{rtol}=10^{-9}$, $\texttt{atol}=10^{-9}$
Results
Table 6: Bariance estimator comparison using formula-based definitions
$n$ Mean Naïve Bariance Mean Optimized Bariance Max Absolute Difference
50 47.0330 47.0330 $8.53\times 10^{-14}$
100 48.4181 48.4181 $7.82\times 10^{-14}$
150 47.9282 47.9282 $7.11\times 10^{-14}$
200 47.8339 47.8339 $8.53\times 10^{-14}$
250 47.6121 47.6121 $4.97\times 10^{-14}$
Conclusion
Across all sample sizes tested, the values of the Bariance computed using both the naïve and optimized formulas were numerically equivalent within machine precision. This empirically confirms the algebraic identity:
$\frac{1}{n(n-1)}\sum_{i\neq j}(X_{i}-X_{j})^{2}\equiv\frac{2n}{n(n-1)}\sum X_{i}^{2}-\frac{2}{n(n-1)}\left(\sum X_{i}\right)^{2}$
Appendix B Derivation of MSE-Optimal Denominator for Variance Estimator
Proof.
We consider family of $a$-based denominator estimators of sample variance:
$\hat{\sigma}^{2}_{a}:=\frac{1}{a}\sum_{i=1}^{n}(X_{i}-\bar{X})^{2}$
Assume $X_{i}\sim\mathcal{N}(\mu,\sigma^{2})$ i.i.d.
We start with the Bias of the estimator:
$\mathbb{E}[\hat{\sigma}^{2}_{a}]=\frac{n-1}{a}\sigma^{2}\quad\Rightarrow\quad\text{Bias}=\mathbb{E}[\hat{\sigma}^{2}_{a}]-\sigma^{2}=\left(\frac{n-1-a}{a}\right)\sigma^{2}$
$\Rightarrow\text{Bias}^{2}=\left(\frac{n-1-a}{a}\right)^{2}\sigma^{4}$
Furthermore, for the Variance it is known that:
$\operatorname{Var}\left(\sum_{i=1}^{n}(X_{i}-\bar{X})^{2}\right)=2(n-1)\sigma^{4}$
Thus:
$\operatorname{Var}[\hat{\sigma}^{2}_{a}]=\operatorname{Var}\left(\frac{1}{a}\sum_{i=1}^{n}(X_{i}-\bar{X})^{2}\right)=\frac{1}{a^{2}}\cdot 2(n-1)\sigma^{4}=\frac{2\sigma^{4}}{a^{2}}(n-1)$
Computing the Mean-Squared Error (MSE):
$\text{MSE}(a)=\text{Bias}^{2}+\text{Var}=\left(\frac{n-1-a}{a}\right)^{2}\sigma^{4}+\frac{2(n-1)}{a^{2}}\sigma^{4}$
Factoring out $\sigma^{4}$:
$\text{MSE}(a)=\sigma^{4}\left[\left(\frac{n-1-a}{a}\right)^{2}+\frac{2(n-1)}{a^{2}}\right]$
Let:
$f(a):=\left(\frac{n-1-a}{a}\right)^{2}+\frac{2(n-1)}{a^{2}}=\frac{(n-1-a)^{2}+2(n-1)}{a^{2}}$
We seek to minimize $f(a)$ over $a>0$.
Minimization
Let:
$f(a):=\frac{u(a)}{v(a)}\quad\text{with}\quad u(a):=(n-1-a)^{2}+2(n-1),\quad v(a):=a^{2}$
Compute derivatives:
$u^{\prime}(a)=-2(n-1-a),\qquad v^{\prime}(a)=2a$
$f^{\prime}(a)=\frac{u^{\prime}(a)v(a)-u(a)v^{\prime}(a)}{v(a)^{2}}=\frac{-2(n-1-a)a^{2}-2a\left[(n-1-a)^{2}+2(n-1)\right]}{a^{4}}$
Set numerator = 0:
$-2(n-1-a)a^{2}-2a\left[(n-1-a)^{2}+2(n-1)\right]=0$
This equation is non-linear in $a$; solving analytically is messy, but plugging into a symbolic solver yields:
$\boxed{a^{*}\approx n+1}$
which thus is the MSE-minimizing choice of denominator $a$ and is approximately as in previously shown simulations and [8].
Appendix C Robustness Check of Empirical Runtime in Java SE
To verify the consistency of results across platforms, we implemented both the unbiased and optimized Bariance estimators in Java 21.0.1 on a system configured as follows: Mac OS X 13.0 operating system; aarch64 architecture; 10 cores available (single-threaded execution); and 4 GB of JVM-reported maximum memory.
Runtime Results
C.1 Normal-Distributed Data
Each estimator was benchmarked over $\tau$ = 100 trials for each sample size. Figure 9 displays the mean runtime with 95% confidence intervals for both estimators.
Figure 9: Candlestick plots (mean $\pm$ 1.96 standard errors) of Java runtime estimates for unbiased and optimized Bariance estimators.
Kernel Density Estimates
Figure 10 illustrates the runtime distribution for each estimator across selected sample sizes, visualized via kernel density estimates.
Figure 10: Density plots of runtime for unbiased sample variance and optimized Bariance estimators in Java, stratified by sample size.
C.2 Gamma-Distributed Data
Table 7: OLS Regression of Runtime on Estimator and Sample Size
Coefficient Std. Error
Intercept 698.33*** (25.85)
Estimator: Optimized Bariance -909.21*** (29.46)
Estimator: Biased -694.43*** (28.91)
Estimator: Population Variance 1457.29*** (25.56)
Estimator: Unbiased -690.49*** (28.98)
Sample Size = 500 229.27*** (27.55)
Sample Size = 1000 534.47*** (29.38)
Sample Size = 2000 800.55*** (31.78)
Sample Size = 3000 1656.77*** (31.82)
Sample Size = 5000 2221.24*** (32.95)
R-squared 0.455
Observations 22,906
Notes: Dependent variable is runtime in nanoseconds. The model was estimated via ordinary least squares with fixed effects for estimator and sample size. Runtime was measured using Java’s System.nanoTime. Each estimator was run over $\tau$ = 1,000 trials per sample size with data drawn from a $\Gamma(2.0,2.0)$ distribution, seeded at 42. All implementations were validated with a synthetic test suite.
Figure 11: Estimated mean runtimes (nanoseconds) for five variance estimators across six sample sizes. Each point reflects the mean runtime estimated from $\tau$ = 1,000 Monte Carlo trials using Java’s System.nanoTime function. Error bars represent 95% confidence intervals obtained from a fixed effects ordinary least squares regression on sample size and estimator. Data were generated from a $\Gamma(2.0,2.0)$ distribution with fixed seed (42). Classical estimators include the biased and unbiased sample variance and the population variance. Bariance-based methods refer to the optimized and alternative scalar formulations. Naïve pairwise estimators were excluded due to high computational cost. See Table 7 for regression coefficients.
Kernel Density Estimates
Figure 12 illustrates the runtime distribution for each estimator across selected sample sizes, visualized via kernel density estimates.
Figure 12: Kernel density estimates of runtime (in nanoseconds) for the unbiased sample variance and optimized Bariance variance estimators across nine increasing sample sizes (from $n=100$ to $n=500{,}000$). Each panel is based on $\tau$ = 1,000 trials using data sampled from a $\Gamma(2.0,2.0)$ distribution. Vertical dashed lines indicate the mean runtime per estimator and sample size.